# TRABSA: Interpretable Sentiment Analysis of Tweets using Attention-based BiLSTM and Twitter-RoBERTa

### v1

@@ 3 Methodology
Our proposed methodological framework outlines a structured approach to SA, encompassing several key stages to ensure robustness and effectiveness in model development and evaluation, as shown in Figure 1. The first stage involves gathering relevant data for SA. We extend existing datasets by collecting additional tweets from diverse sources to enhance the dataset’s representativeness and coverage. Following data collection, we perform cleaning and preprocessing tasks to ensure the quality and consistency of the data. This includes removing noise, expanding contractions, handling duplicates, emojis, and missing tweets, and standardizing text formats. The next step involves labeling the data into positive, neutral, and negative sentiments. We leverage the latest updated RoBERTa-based pre-trained transformer model for tokenization and sentiment labeling, enabling accurate and efficient text data processing. The dataset is divided into training, validation, and test sets after it has been labeled. This enables us to use the test set to assess the model’s performance on untested data, refine hyperparameters using the validation set, and train the model on a portion of the data. In this step, we build the SA model based on our suggested hybrid DL architecture. Following model creation, we compare the trained model’s performance against baseline models or current state-of-the-art methods. We examine the model’s ability to correctly classify sentiments using a variety of metrics, including accuracy, precision, recall, and F1-score. Finally, we employ XAI techniques to interpret the model’s predictions and gain insights into its decision-making process. This involves analyzing the model’s internal mechanisms, such as attention weights or feature importance, to understand the factors influencing its predictions and enhance model interpretability.
3.1 Data Collection and Preprocessing
@@
In this section, we provide a comprehensive overview of the data collection and preprocessing procedures undertaken in our study, which laid the foundation for robust SA of tweets.
3.1.1 Data Sources
@@
Benchmark Dataset: The benchmark dataset, which forms the basis of our research, was first assembled and curated by [Qi & Shabrina, 2023]. It functions as a standard by which our proposed model’s performance is measured. Interestingly, our model outperformed this benchmark dataset, indicating significant progress in SA. This reference dataset consists of tweets with geotags that were sent from well-known cities in the United Kingdom during the third nationwide COVID-19 shutdown. Figure 2 shows the tweets gathered from the three stages in the UK. This group of cities includes Greater London, Bristol, South Hampton, Birmingham, Manchester, Liverpool, Newcastle, Leeds, Sheffield, and Nottingham. Over the course of three weeks, from January 6, 2021, to July 18, 2021, 77,332 tweets were gathered. 29,923 tweets were gathered in the first stage, 24,689 in the second, and 22,720 in the third. Major cities such as London, Manchester, Birmingham, and Liverpool were the source of the majority of tweets, with London having the highest count with 37,678. Smaller cities, like Newcastle, had just 852 tweets in a six-month period. The data distribution is in phases, with the first stage having the greatest data and the third stage having the least, as Figure 2 illustrates. While Newcastle’s contribution was connected with its population and density, London consistently supplied the most data.
@@
Extended Datasets: To augment our research’s cross-cultural dimension and overcome the geographic limitations of the benchmark dataset proposed by Qi and Shabrina, we extended the existing UK COVID-19 dataset [Qi & Shabrina, 2023]. The tweets were sourced through a combination of data extraction tools, specifically Twint and the Twitter Academic API. These tools were chosen because of their ability to acquire tweets with geolocation information, which is crucial for conducting geographical analyses. However, it should be noted that only a few 1% of Twitter users actively opt to share their geographic location when composing tweets, and this feature is not enabled by default [Sloan & Morgan, 2015]. The extended datasets, comprising the Global Twitter COVID-19 Dataset and the USA Twitter COVID-19 Dataset, are publicly available in the Extended Covid Twitter Datasets repository [Jahin, 2023].
To ensure a comprehensive dataset, we merged the data collected by Twint and the Twitter Academic API. This amalgamation allowed us to access a larger volume of tweets. In identifying tweets related to the COVID-19 pandemic, we employed specific keywords such as “corona” or “covid” in the Twint search configurations and the query field of the Twitter Academic API. This search strategy enabled us to extract tweets and associated hashtags containing these pertinent terms.
@@
    Extended Global COVID-19 Dataset: This extension involved the comprehensive scraping of 411,885 tweets from 32 English-speaking countries. This dataset expansion allowed us to capture sentiment variations across diverse English-speaking regions, as illustrated in Figure 3. In particular, cities such as the “United States,” the “United Kingdom,” “Australia,” and “New Zealand” exhibit high tweet volumes, while several other cities have comparatively lower tweet counts. Figures 4 and 5 illustrate word clouds and word frequencies within tweets of the extended datasets. Figures 4 (left) and 5 (left) represent a visual summary of the most frequently occurring words in a vast dataset related to the COVID-19 pandemic. At the center of this cloud is the word “covid,” which dominates with a staggering 226,463 mentions. Other significant terms around it, such as “vaccine,” “case,” “test,” and “people,” are indicative of the key topics and concerns worldwide during the pandemic. Words like “death,” “pandemic,” and “health” also hold prominence, highlighting the gravity of public health issues. Additionally, terms like “Trump” and “government” suggest the political dimensions entwined with the pandemic discourse.
@@
    Extended USA COVID-19 Dataset: In addition to the international extension, we further enriched our data by creating an extended dataset focusing exclusively on the United States. This dataset comprised 7,500 tweets meticulously scraped from U.S.-based sources. Including this dataset allows for a closer examination of sentiment dynamics within a specific geographical context. Figures 4 (right) and 5 (right) specific to the United States reveal notable trends and sentiments within the country during the COVID-19 pandemic. In this dataset, words like “corona,” “coronavirus,” and “covid” are prominent, underlining the ubiquitous presence of these terms in American discussions. Interestingly, words like “fool” and “joke” appear, possibly reflecting a spectrum of attitudes towards the pandemic response. Negative expressions like “shit,” “fuck,” and “die” also emerge, suggesting the emotional intensity and frustration associated with the situation. Terms like “test” and “case” point towards testing and infection rates concerns, while “april” indicates a temporal reference.
Figure 4: This word cloud visualizes the most frequently occurring words in the “Global” (left) and “Only USA” (right) datasets of COVID-19-related tweets. The size of each word corresponds to its frequency in the dataset.
Figure 5: This line plot displays the word frequency within tweets in the “Global” (left) and “Only USA” (right) COVID-19 datasets. It provides insights into the most commonly used words regarding COVID-19 discussions among Twitter users.
External Datasets: In order to evaluate the robustness and generalizability of our proposed model, we incorporated four external datasets sourced from Kaggle, including the Twitter Dataset, Reddit Dataset, Apple Twitter Dataset, and US Airline Twitter Dataset.
@@
3.1.2 Data Cleaning and Preprocessing
As the previous literature indicated, preprocessing the raw Twitter data was essential to guarantee the accuracy and reliability of our SA because of their informal and unstructured character [Naseem et al., 2021]. Our thorough data-cleaning procedure included the following crucial steps:
@@ 3.1.2 Data Cleaning and Preprocessing
    Removal of Irrelevant Elements: We methodically eliminated any superfluous content that has no bearing on SA, including hashtags (#subject), stated usernames (@username), and any hyperlinks beginning with “www,” “http,” or “https.” We also removed terms that were less than two characters and stop words. Stop words, though common in text, often lack significant sentiment polarity. Despite being often used in texts, stop words frequently lack strong emotive polarity. It’s important to note that negations like “not” and “no” were kept in because removing them may change the sense of whole sentences.
@@
    Handling Repeated Characters: Some users utilize repeating characters in their tweets to highlight intense feelings. Words not present in standard lexicons were transformed into their correct forms to standardize such expressions. For instance, “sooooo goooood” was normalized to “so good.”
@@
    Extending Contractions: Removing punctuation after a contraction, such “isn’t” or “don’t,” presented difficulties. They were expanded into their full forms to maintain the meaningfulness of contractions. For instance, “isn’t” became “is not.”
@@
3.1.3 Word Embeddings
We used a variety of word embedding approaches in this work to extract contextual and semantic information from our textual material. Our NLP tasks performed much better thanks to these embeddings. The word embedding techniques we used in our studies are summarized in the next subsections.
@@ 3.1.3 Word Embeddings
Term Frequency-Inverse Document Frequency (TF-IDF): We utilized the TF-IDF embeddings by employing the ‘TfidfVectorizer‘ function from the scikit-learn module, which allocates weights to words according to their significance in individual tweets and their scarcity throughout the complete dataset. Additionally, it made it easier to down-weight frequent keywords, which allowed our models to concentrate on words that were more informative.
Word2Vec: One of Word2Vec’s advantages is that it can record semantic similarities between words, which makes text data analysis more sophisticated. It represents words as dense vectors in a continuous vector space by utilizing neural networks. The ‘word_tokenize‘ function from the NLTK library was utilized to tokenize our tweets, as it allows for the breakdown of sentences into individual words. Using the Gensim package, a Word2Vec model was produced with the vector size, window size, and skip-gram model set. By using a continuous vector space, this approach was able to express words as vectors. To encode full tweets as vectors, two methods were used: sum vectorization and average vectorization.
Pre-trained Transformers: In our research, we harnessed the power of pre-trained transformer-based models from the Hugging Face Transformers library to leverage contextual embeddings for text data. Three distinct transformer models were employed, each bringing unique capabilities to the analysis. The ‘distilbert-base-uncased’ model, known for its efficiency and lightweight nature, was selected for its suitability in scenarios where computational resources are constrained. It produces context-aware word embeddings considering each word’s left and right context. We used a state-of-the-art ‘cardiffnlp/twitter-roberta-base-sentiment-latest’ model, updated in 2022, to capture sentiment-specific nuances in a tweet. This model was trained on an extensive dataset of approximately 124 million tweets collected from January 2018 to December 2021. This model was designed for English text and was a robust foundation for our SA endeavors. Additionally, we incorporated ‘sentence-transformers/all-MiniLM-L6-v2,’ a sophisticated tool that transforms tweets into a dense vector space of 384 dimensions. It transforms entire sentences into fixed-dimensional vectors while maintaining semantic information.
@@
3.2 Unsupervised Text Labeling
In ML, manually labeling vast amounts of text data manually can be time-consuming. To address this challenge and expedite the labeling process, we leveraged lexicon-based methods, specifically TextBlob, VADER, and SentiWordNet, to automatically assign sentiment scores to tweets. Our sentiment classification scheme employed three categories: positive (assigned a value of 1), negative (assigned -1), and neutral (assigned 0).
@@ 3.2 Unsupervised Text Labeling
3.3 Models Used
3.3.1 Traditional ML Models
Our analysis encompasses a diverse set of models traditional ML models employed in our study, including traditional base models, their stacked ensembles, and voting classifiers. The objective was to thoroughly assess these models’ performance using various text representations. The models utilized were as follows:
@@ 3.3.1 Traditional ML Models
    Random Forest (RF): RF is an ensemble learning technique combining many decision tree predictions. It is renowned for being resilient and capable of managing high-dimensional data.
@@
    Naive Bayes (NB): NB is a probabilistic classifier based on the Bayes theorem and supposing feature independence. It works especially effectively for jobs involving text categorization.
@@
    Support Vector Machine (SVM): SVM is an effective classifier that looks for the optimal hyperplane in a high-dimensional space to divide data points. It works well for classification that is both linear and non-linear.
@@
    Gradient Boosting Machine (GBM): Using an ensemble learning approach called GBM, decision trees are constructed one after the other with an emphasis on the errors made by the earlier trees. It frequently results in excellent prediction performance.
@@
    LightGBM (LGBM): LightGBM is a gradient-boosting framework designed for efficiency and speed. It uses a histogram-based approach for tree construction.
@@
3.3.2 Deep Neural Networks (DNNs)
@@
We introduced the ‘BiLSTM + 3 Hidden Dense Layers’ model to explore the nuances of DL further. This architecture incorporated a BiLSTM layer, enabling the network to capture sequential dependencies in the input data. Following the BiLSTM layer, three additional dense layers (512, 256, and 128 units) were used to distill the features further. Dropout was applied to enhance model generalization.
@@
Our systematic exploration of these DNN architectures, coupled with a comprehensive array of text representations, provided valuable insights into their performance and allowed us to tailor our approach to the intricacies of tweet SA. This diversity of models served as a critical asset in our research, enabling us to select the most effective strategies for this challenging task.
3.3.3 Proposed TRABSA Model
The proposed TRABSA model presents a systematic and effective architecture for SA, as shown in Figure 6. The model utilizes the ‘cardiffnlp/twitter-roberta-base-sentiment-latest’ pre-trained transformer, capitalizing on its contextual understanding of the text. The architecture begins with input layers, including ‘input_ids’ and ‘attention_mask,’ where a maximum sequence length of 256 tokens is utilized. RoBERTa is used for its excellent performance in tasks involving NLP. It encodes the input tweet text and generates contextual embeddings.
@@ 3.3.3 Proposed TRABSA Model
Figure 8 comprehensively evaluates the TRABSA model’s performance in tweet SA. The top left corner showcases the model’s classification metrics plot, illustrating precision, recall, and F1-score metrics for each sentiment class. Moving clockwise, the confusion matrix provides a detailed analysis of the model’s classification performance by comparing predicted sentiment labels with actual labels. The loss vs. epoch curve illustrates its training and validation loss over successive epochs, while the training and validation accuracy vs. epoch curve depicts the model’s learning progress and convergence. Together, these visualizations offer insights into the TRABSA model’s classification accuracy, convergence, and optimization process, aiding researchers in assessing its performance and identifying areas for improvement.
The TRABSA model’s architecture combines the strengths of RoBERTa’s contextual embeddings, BiLSTM’s sequence modeling, attention mechanisms for capturing interdependencies, and a well-designed set of dense layers to improve SA accuracy for tweet data. This architecture demonstrated superior performance compared to other models, making it a noteworthy addition to the field of SA.
4 Experimental Analysis and Results
4.1 Experimental Setup

### v2

@@ Introduction
Our proposed methodological framework outlines a structured approach to SA, encompassing several key stages to ensure robustness and effectiveness in model development and evaluation, as shown in Figure 1. The first stage involves gathering relevant data for SA. We extend existing datasets by collecting additional tweets from diverse sources to enhance the dataset’s representativeness and coverage. Following data collection, we perform cleaning and preprocessing tasks to ensure the quality and consistency of the data. This includes removing noise, expanding contractions, handling duplicates, emojis, and missing tweets, and standardizing text formats. The next step involves labeling the data into positive, neutral, and negative sentiments. We leverage the latest updated RoBERTa-based pre-trained transformer model for tokenization and sentiment labeling, enabling accurate and efficient text data processing. The dataset is divided into training, validation, and test sets after it has been labeled. This enables us to use the test set to assess the model’s performance on untested data, refine hyperparameters using the validation set, and train the model on a portion of the data. In this step, we build the SA model based on our suggested hybrid DL architecture. Following model development, we compare the trained model’s performance against baseline models or current state-of-the-art methods. We examine the model’s ability to correctly classify sentiments into positive, neutral, and negative categories using a variety of metrics, including accuracy, precision, recall, and F1-score. Finally, we employ XAI techniques to interpret the model’s predictions and gain insights into its decision-making process. This involves analyzing the model’s internal mechanisms, such as attention weights or feature importance, to understand the factors influencing its predictions and enhance model interpretability.
Data Collection and Preprocessing
@@
In this section, we provide a comprehensive overview of the data collection and preprocessing procedures undertaken in our study, which laid the foundation for robust sentiment analysis of tweets.
Data Sources
@@
Benchmark Dataset: The benchmark dataset, which forms the basis of our research, was first assembled and curated by [16]. It functions as a standard by which our proposed model’s performance is measured. Interestingly, our model outperformed this benchmark dataset, indicating significant progress in SA. This reference dataset consists of tweets with geotags from well-known cities in the United Kingdom during the third nationwide COVID-19 shutdown. Figure 2 shows the tweets gathered from the three stages in the UK. This group of cities includes Greater London, Bristol, South Hampton, Birmingham, Manchester, Liverpool, Newcastle, Leeds, Sheffield, and Nottingham. Over the course of three weeks, from January 6, 2021, to July 18, 2021, 77,332 tweets were gathered. 29,923 tweets were gathered in the first stage, 24,689 in the second, and 22,720 in the third. Major cities such as London, Manchester, Birmingham, and Liverpool were the source of most tweets, with London having the highest count with 37,678. Smaller cities, like Newcastle, had just 852 tweets in a six-month period. The data distribution is in phases, with the first stage having the greatest data and the third stage having the least, as Figure 2 illustrates. While Newcastle’s contribution was connected with its population and density, London consistently supplied the most data.
@@
Extended Datasets: To augment our research’s cross-cultural dimension and overcome the geographic limitations of the benchmark dataset proposed by Qi and Shabrina, we extended the existing UK COVID-19 dataset [16]. The tweets were sourced through a combination of data extraction tools, specifically Twint and the Twitter Academic API. These tools were chosen because of their ability to acquire tweets with geolocation information, which is crucial for conducting geographical analyses. However, it should be noted that only a few 1% of Twitter users actively opt to share their geographic location when composing tweets, and this feature is not enabled by default [40]. The extended datasets, comprising the Global Twitter COVID-19 Dataset and the USA Twitter COVID-19 Dataset, are publicly available in the Extended Covid Twitter Datasets repository [41].
To ensure a comprehensive dataset, we merged the data collected by Twint and the Twitter Academic API. This amalgamation allowed us to access a larger volume of tweets. In identifying tweets related to the COVID-19 pandemic, we employed specific keywords such as “corona" or “covid" in the Twint search configurations and the query field of the Twitter Academic API. This search strategy enabled us to extract tweets and associated hashtags containing these pertinent terms.
@@
    Extended Global COVID-19 Dataset: This extension involved the comprehensive scraping of 411,885 tweets from 32 English-speaking countries. This dataset expansion allowed us to capture sentiment variations across diverse English-speaking regions, as illustrated in Figure 3. In particular, cities such as the “United States," the “United Kingdom," “Australia," and “New Zealand" exhibit high tweet volumes, while several other cities have comparatively lower tweet counts. Figure 4 illustrates word clouds and word frequencies within tweets of the extended datasets. Figure 4 (left) represents a visual summary of the most frequently occurring words in a vast dataset related to the COVID-19 pandemic. At the center of this cloud is the word “covid," which dominates with a staggering 226,463 mentions. Other significant terms around it, such as “vaccine," “case," “test," and “people," indicate the key topics and concerns worldwide during the pandemic. Words like “death," “pandemic," and “health" also hold prominence, highlighting the gravity of public health issues. Additionally, terms like “Trump" and “government" suggest the political dimensions entwined with the pandemic discourse.
@@
    Extended USA COVID-19 Dataset: In addition to the international extension, we further enriched our data by creating an extended dataset focusing exclusively on the United States. This dataset comprised 7,500 tweets meticulously scraped from U.S.-based sources. Including this dataset allows for a closer examination of sentiment dynamics within a specific geographical context. Figure 4 (right) specific to the United States reveals notable trends and sentiments within the country during the COVID-19 pandemic. In this dataset, words like “corona," “coronavirus," and “covid" are prominent, underlining the ubiquitous presence of these terms in American discussions. Interestingly, words like “fool" and “joke" appear, possibly reflecting a spectrum of attitudes towards the pandemic response. Negative expressions like “shit," “fuck," and “die" also emerge, suggesting the emotional intensity and frustration associated with the situation. Terms like “test" and “case" point towards testing and infection rates concerns, while “april" indicates a temporal reference.
Figure 4: This word cloud visualizes the most frequently occurring words in the “Global" (left) and “Only USA" (right) datasets of COVID-19-related tweets. The size of each word corresponds to its frequency in the dataset.
External Datasets: In order to evaluate the robustness and generalizability of our proposed model, we incorporated four external datasets sourced from Kaggle, including the Twitter and Reddit Dataset, Apple Dataset, and US Airline Dataset (see Data Availability section).
@@
Data Cleaning and Preprocessing
As a previous study has indicated, preprocessing the raw Twitter data was essential to guarantee the accuracy and reliability of our SA because of their informal and unstructured character [42]. Our thorough data-cleaning procedure included the following crucial steps:
@@
    Removal of Irrelevant Elements: We methodically eliminated any superfluous content that has no bearing on SA, including hashtags (#subject), stated usernames (@username), and any hyperlinks beginning with “www," “http," or “https." We also removed terms that were less than two characters and stop words. Stop words, though common in text, often lack significant sentiment polarity. Despite being often used in texts, stop words frequently lack strong emotive polarity. It’s important to note that negations like “not" and “no" were kept in because removing them may change the sense of whole sentences.
@@
    Handling Repeated Characters: Some users utilize repeating characters in their tweets to highlight intense feelings. Words not present in standard lexicons were transformed into their correct forms to standardize such expressions. For instance, “sooooo goooood" was normalized to “so good."
@@
    Extending Contractions: Removing punctuation after a contraction, such “isn’t" or “don’t," presented difficulties. They were expanded into their full forms to maintain the meaningfulness of contractions. For instance, “isn’t" became “is not."
@@
Word Embeddings
We used various word embedding approaches in this work to extract contextual and semantic information from our textual material. Our NLP tasks performed much better thanks to these embeddings. The word embedding techniques we used in our studies are summarized in the next subsections.
@@
Term Frequency-Inverse Document Frequency (TF-IDF): We utilized the TF-IDF embeddings by employing the ‘TfidfVectorizer‘ function from the scikit-learn module, which allocates weights to words according to their significance in individual tweets and their scarcity throughout the complete dataset. Additionally, it made it easier to down-weight frequent keywords, which allowed our models to concentrate on more informative words.
Word2Vec: One of Word2Vec’s advantages is that it can record semantic similarities between words, which makes text data analysis more sophisticated. Using neural networks, it represents words as dense vectors in a continuous vector space. The ‘word_tokenize‘ function from the NLTK library was utilized to tokenize our tweets, as it allows for the breakdown of sentences into individual words. Using the Gensim package, a Word2Vec model was produced with the vector size, window size, and skip-gram model set. By using a continuous vector space, this approach was able to express words as vectors. Two methods were used to encode full tweets as vectors: sum vectorization and average vectorization.
Pre-trained Transformers: In our research, we harnessed the power of pre-trained transformer-based models from the Hugging Face Transformers library to leverage contextual embeddings for text data. Three distinct transformer models were employed, each bringing unique capabilities to the analysis. The ‘distilbert-base-uncased’ model, known for its efficiency and lightweight nature, was selected for its suitability in scenarios where computational resources are constrained. It produces context-aware word embeddings that consider each word’s left and right context. We used a state-of-the-art ‘cardiffnlp/twitter-roberta-base-sentiment-latest’ model, updated in 2022, to capture sentiment-specific nuances in a tweet. This model was trained on an extensive dataset of approximately 124 million tweets collected from January 2018 to December 2021. This model was designed for English text and was a robust foundation for our SA endeavors. Additionally, we incorporated ‘sentence-transformers/all-MiniLM-L6-v2,’ a sophisticated tool that transforms tweets into a dense vector space of 384 dimensions. It transforms entire sentences into fixed-dimensional vectors while maintaining semantic information.
@@
Unsupervised Text Labeling
In ML, labeling vast amounts of text data manually can be time-consuming. To address this challenge and expedite the labeling process, we leveraged lexicon-based methods, specifically TextBlob, VADER, and SentiWordNet, to automatically assign sentiment scores to tweets. Our sentiment classification scheme employed three categories: positive (assigned a value of 1), negative (assigned -1), and neutral (assigned 0).
@@
Models Used
Traditional ML Models
Our analysis encompasses a diverse set of models traditional ML models, including traditional base models, their stacked ensembles, and voting classifiers. The aim was to comprehensively evaluate the performance of these models using different text representations. The following models were used:
@@
    Random Forest (RF): RF is an ensemble learning method that aggregates the predictions of multiple decision trees. It is known for its robustness and ability to handle high-dimensional data.
@@
    Naive Bayes (NB): NB is a probabilistic classifier based on Bayes’ theorem, assuming independence among features. It is particularly well-suited for text classification tasks.
@@
    Support Vector Machine (SVM): SVM is a powerful classifier that aims to find a hyperplane that best separates data points in a high-dimensional space. It is effective for both linear and non-linear classification.
@@
    Gradient Boosting Machine (GBM): GBM is an ensemble learning technique that builds decision trees sequentially, focusing on the mistakes of the previous trees. It often leads to strong predictive performance.
@@
    LightGBM (LGBM): LightGBM is a gradient-boosting framework for efficiency and speed. It uses a histogram-based approach for tree construction.
@@
Deep Neural Networks (DNNs)
@@
We introduced the ‘BiLSTM + 3 Hidden Dense Layers’ model to explore the nuances of DL further. This architecture incorporated a bidirectional Long Short-Term Memory (BiLSTM) layer, enabling the network to capture sequential dependencies in the input data. Following the BiLSTM layer, three additional dense layers (512, 256, and 128 units) were used to distill the features further. Dropout was applied to enhance model generalization.
@@
Our systematic exploration of these DNN architectures, coupled with a comprehensive array of text representations, provided valuable insights into their performance and allowed us to tailor our approach to the intricacies of tweet sentiment analysis. This diversity of models served as a critical asset in our research, enabling us to select the most effective strategies for this challenging task.
Proposed TRABSA Model
The proposed TRABSA model presents a systematic and effective architecture for sentiment analysis, as shown in Figure 5. The model utilizes the ‘cardiffnlp/twitter-roberta-base-sentiment-latest’ pre-trained transformer, capitalizing on its contextual understanding of the text. The architecture begins with input layers, including ‘input_ids’ and ‘attention_mask,’ where a maximum sequence length of 256 tokens is utilized. RoBERTa is used for its excellent performance in tasks involving natural language understanding. It encodes the input tweet text and generates contextual embeddings.
@@
Figure 7 comprehensively evaluates the TRABSA model’s performance in tweet sentiment analysis. The top left corner showcases the model’s classification metrics plot, illustrating precision, recall, and F1-score metrics for each sentiment class. Moving clockwise, the confusion matrix provides a detailed analysis of the model’s classification performance by comparing predicted sentiment labels with actual labels. The loss vs. epoch curve illustrates its training and validation loss over successive epochs, while the training and validation accuracy vs. epoch curve depicts the model’s learning progress and convergence. Together, these visualizations offer insights into the TRABSA model’s classification accuracy, convergence, and optimization process, aiding researchers in assessing its performance and identifying areas for improvement.
The TRABSA model’s architecture combines the strengths of RoBERTa’s contextual embeddings, BiLSTM’s sequence modeling, attention mechanisms for capturing interdependencies, and a well-designed set of dense layers to improve sentiment analysis accuracy for tweet data. This architecture demonstrated superior performance compared to other models, making it a noteworthy addition to the field of sentiment analysis.
Results
Experimental Setup